OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9.

OpenStudy (anonymous):

slope = 4/5 and y-intercept = −9/5 slope = −4/5 and y-intercept = 9/5 slope = −4/5 and y-intercept = −9/5 slope = 4/5 and y-intercept = 9/5

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

C